OpenStudy (christos):

y = mx - c I want you to tell me why y = x1 if x1=x2 ?

hero (hero):

um....what?

OpenStudy (anonymous):

What? That isn't necessarily true.

OpenStudy (anonymous):

y = mx - c y = x1 x1=mx-c x2=mx-c lol that is pretty much all the algebraic manipulation you can do.